OpenStudy (swaqqout_kid1):

Shannon wants to buy cupcakes for her friend's birthday party. The bakery she goes to sells cupcakes for $3.25 each. The bakery also sells gift cards for $2.00, and Shannon decides to buy one card. If Shannon buys 7 cupcakes, how much money does she spend at the bakery? @sangya21

OpenStudy (anonymous):

3.25(7) + 2.00

sammixboo (sammixboo):

Okie dokie! So we have \(\rm (3.25c)~ + ~(2.00g)\) C = Number of Cupcakes = 7 G = Number of Gift Cards = 1 So plug in the values :) \(\rm (3.25\times7)~ + ~(2.00\times 1)\) Now multiply both \(\rm 3.25\times 7\) and \(\rm2.00 \times 1\) and once you have found the product of both of the multiplication problems, you add them together and that is your answer!

sammixboo (sammixboo):

Sorry wrong post ;(

sammixboo (sammixboo):

If you need anymore help tag me or Sangya @sammixboo or @sangya21 :)